OpenStudy (anonymous):

Prove that the given equation is an identity. (sin θ−cos θ)2 = 1−sin 2θ

OpenStudy (anonymous):

\[sin(2\theta) = 2sin(\theta)cos(\theta)\]

OpenStudy (zarkon):

use \sin for a nicer looking sine function :) \[\sin(\theta)\]

OpenStudy (zarkon):

also use \[\sin^2(\theta)+\cos^2(\theta)=1\]